OpenStudy (anonymous):

Which equation represents the line passing through the points (1, 1) and (−2, 13)?

myininaya (myininaya):

Do you know what an equation of a line looks like?

OpenStudy (anonymous):

The Formula you should use is...y2-y1/x2-x1=y-y2/x - x2...x1,y1=1,1 and x2,y2=-2/13..put this ya will get it

OpenStudy (anonymous):

Explain?

OpenStudy (anonymous):

Do you know how to find the slopes of given two points?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

suppose the two points are x1,y1 and x2,y2....the slope is given as m=y2-y1/x2-x1.....okay now anypoint on that line (x,y) also have the same slopes with either of those points....y-y1/x-x1=m=y2-y1/x2-x1

OpenStudy (anonymous):

im having a hard time following what you're saying

OpenStudy (anonymous):

ya know abot slope right?